# Enhancing Dialogue Annotation with Speaker Characteristics Leveraging a Frozen LLM

### v1

@@ Abstract
Index Terms:
Large Language Models, Speaker Characterization, Automatic Speaker Verification, Emotion Recognition, Connectors
I Introduction
@@
Section II surveys related work and highlights key differences with existing approaches, which serve as our baselines. Section III describes the datasets used, the architecture of our models, and the experimental protocols, including task definitions and evaluation metrics. In Section IV, we report the performance of our method across selected tasks, before interpreting the results and discussing broader implications in Section V.
II Related Work
@@ II-A Accoustic Encoders
II-B Audio understanding with LLM
@@ III-A Datasets
III-A1 Automatic Speaker Verification
@@
III-A2 Speaker Age and Gender Classification
@@
III-A3 Speech Emotion Recognition
@@
III-A4 Automatic Speech Recognition
@@ III-C Models
III-C1 Speaker Attribute Pipeline
@@
III-C2 Speaker Verification Pipeline
@@
III-D Experiments
Figure 1: Schematic representing the Speaker Attribute tasks performed. The baseline [13] we are comparing to is the frozen Whisper model with a trainable TLTR connector and a LLAMA fine-tuned using LoRa adaptors.
III-D1 Speaker Attributes Extraction
@@
Model outputs are evaluated on their respective test sets using standard metrics, as defined in Section III-B.
@@
Figure 2: Schematic representing the process to simulate speaker verification within a conversation by taking utterances from the same speaker but from different sessions.
TABLE I: Performances on various downstream tasks using different encoders, compared to a baseline using only audio inputs. The best value is bolded, and the second best is in italics. When different encoders of classifiers are used for different tasks, the number of parameters is shown as $NumberOfSystems\times Parameters$
@@ IV Results
IV-A Speaker Attibutes Results
IV-A1 Task-specific connectors
Table I presents the performance of our proposed models across four downstream tasks: gender classification, age prediction, emotion recognition, and automatic speech recognition (ASR). We report results for models based on WavLM and Whisper embeddings, with lightweight task-specific connectors, and compare them against state-of-the-art baselines such as LTU-AS [13], WavLM+MLP [24], and Vox-Profile [25]. The table also includes the number of trainable parameters per model variant.
@@ IV-A2 Universal connector
IV-A3 Comparison with the transcripts
The comparative results obtained using the whisper transcripts are shown in Line 7 of Table I. The ASR task shows a non-negligeable degradation of the WER, from 1.4% in the whisper transcripts relative to the ground truth, to 4.04% on the outputs of the LLAMA Vicunas. The slight degradation observed does not explain entirely the high WER observed using the WavLM and Whisper embeddings mean pooled, which supports the idea of using parallel transcripts in addition to audio embeddings for future dialogue annotations. As expected, age, gender and emotion predictions are significantly worse than previous experiments, with notably emotions being significantly worse than random (10.29% accuracy for 4 classes) and gender being almost random, knowing that 73% of the utterances in the set are Male. The F1-score for gender classification is measured 0.08.
IV-B Speaker Verification Results
TABLE II: Performances on the speaker verification task, considering various numbers of embeddings for each test speaker. We always use one enrollment embedding from a speaker $i$ to compare to $N_{j}+N_{k}$ test embeddings. $N_{j}\geq 1$ is the number of embeddings used for the target speaker, while $N_{k}\geq 0$ is the number of embeddings used for a third speaker $k$, different than both $i$ and $j$.
@@ IV-B Speaker Verification Results
Table II presents our speaker verification experiments results. Line 1 shows the standard one-to-one comparison used for speaker verification, following the couple of enroll/test audio segments of VoxCeleb1-O. On this task, our adapted system yields 12.08% EER, a very high result compared to the original system that yielded 0.80% on the same setup. This first result shows two conclusions:

### v2

@@ Abstract
Keywords: Large Language Models, Speaker Characterization, Automatic Speaker Verification, Emotion Recognition, Connectors
1 Introduction
@@
Section 2 surveys related work and highlights key differences with existing approaches, which serve as our baselines. Section 3 describes the datasets used, the architecture of our models, and the experimental protocols, including task definitions and evaluation metrics. In Section 4, we report the performance of our method across selected tasks, before interpreting the results and discussing broader implications in Section 5.
2 Related Work
@@ 2.1 Accoustic Encoders
2.2 Audio understanding with LLM
@@ 3.1 Datasets
3.1.1 Automatic Speaker Verification
@@
3.1.2 Speaker Age and Gender Classification
@@
3.1.3 Speech Emotion Recognition
@@
3.1.4 Automatic Speech Recognition
@@ 3.3 Models
3.3.1 Speaker Attribute Pipeline
@@
3.3.2 Speaker Verification Pipeline
@@
3.4 Experiments
Fig. 1: Schematic representing the Speaker Attribute tasks performed. The baseline [13] we are comparing to is the frozen Whisper model with a trainable TLTR connector and a LLAMA fine-tuned using LoRa adaptors.
3.4.1 Speaker Attributes Extraction
@@
Model outputs are evaluated on their respective test sets using standard metrics, as defined in Section 3.2.
@@
Fig. 2: Schematic representing the process to simulate speaker verification within a conversation by taking utterances from the same speaker but from different sessions.
Table 1: Performances on various downstream tasks using different encoders, compared to a baseline using only audio inputs. The best value is bolded, and the second best is in italics. When different encoders of classifiers are used for different tasks, the number of parameters is shown as $NumberOfSystems\times Parameters$
@@ 4 Results
4.1 Speaker Attibutes Results
4.1.1 Task-specific connectors
Table 1 presents the performance of our proposed models across four downstream tasks: gender classification, age prediction, emotion recognition, and automatic speech recognition (ASR). We report results for models based on WavLM and Whisper embeddings, with lightweight task-specific connectors, and compare them against state-of-the-art baselines such as LTU-AS [13], WavLM+MLP [24], and Vox-Profile [25]. The table also includes the number of trainable parameters per model variant.
@@ 4.1.2 Universal connector
4.1.3 Comparison with the transcripts
The comparative results obtained using the whisper transcripts are shown in Line 7 of Table 1. The ASR task shows a non-negligeable degradation of the WER, from 1.4% in the whisper transcripts relative to the ground truth, to 4.04% on the outputs of the LLAMA Vicunas. The slight degradation observed does not explain entirely the high WER observed using the WavLM and Whisper embeddings mean pooled, which supports the idea of using parallel transcripts in addition to audio embeddings for future dialogue annotations. As expected, age, gender and emotion predictions are significantly worse than previous experiments, with notably emotions being significantly worse than random (10.29% accuracy for 4 classes) and gender being almost random, knowing that 73% of the utterances in the set are Male. The F1-score for gender classification is measured 0.08.
4.2 Speaker Verification Results
Table 2: Performances on the speaker verification task, considering various numbers of embeddings for each test speaker. We always use one enrollment embedding from a speaker $i$ to compare to $N_{j}+N_{k}$ test embeddings. $N_{j}\geq 1$ is the number of embeddings used for the target speaker, while $N_{k}\geq 0$ is the number of embeddings used for a third speaker $k$, different than both $i$ and $j$.
@@ 4.2 Speaker Verification Results
Table 2 presents our speaker verification experiments results. Line 1 shows the standard one-to-one comparison used for speaker verification, following the couple of enroll/test audio segments of VoxCeleb1-O. On this task, our adapted system yields 12.08% EER, a very high result compared to the original system that yielded 0.80% on the same setup. This first result shows two conclusions: